OpenStudy (kevin):

Find the easiest way to solve 3^2008 (10^2013 + 5^2012 + 2^2011) ------------------------------------ 5^2012 (6^2010 + 3^2009 + 2^2008)

OpenStudy (kevin):

@zzr0ck3r

OpenStudy (kevin):

Don't need to find the final answer, just how is the easiest way to solve it.

OpenStudy (kevin):

@ganeshie8

OpenStudy (kevin):

@princeharryyy

OpenStudy (princeharryyy):

I don't know what are you solving for. If you are considering adding them. Sorry. I don't think you can. But, remainder is possible.

OpenStudy (princeharryyy):

I don't know what are you solving for. If you are considering adding them. Sorry. I don't think you can. But, remainder is possible.

OpenStudy (kevin):

This was the question that my teacher gave to me. You don't need to solve it anyway. You just should find what is 'the easiest way' to solve it.

OpenStudy (skullpatrol):

Note: 10=2*5 and 6=2*3

OpenStudy (kevin):

How about '+' symbols?

OpenStudy (skullpatrol):

You're going to have to use the distributive property too.

OpenStudy (skullpatrol):

$$ 3^{2008} (10^{2013} + 5^{2012} + 2^{2011})= 3^{2008} ((2•5)^{2013} + 5^{2012} + 2^{2011})$$

OpenStudy (kevin):

can we divide numerator with denominator?

OpenStudy (skullpatrol):

hmmm... this is a tough one

OpenStudy (skullpatrol):

$$\huge\dfrac{ 3^{2008} ((2•5)^{2013} + 5^{2012} + 2^{2011})}{5^{2012} ((2•3)^{2010} + 3^{2009} + 2^{2008})}$$ $$\huge = \dfrac{ 3^{2008} (2^{2013}•5^{2013} + 5^{2012} + 2^{2011})}{5^{2012} (2^{2010}•3^{2010} + 3^{2009} + 2^{2008})}$$

OpenStudy (kevin):

I still get no clue.. why if we distribute it like that, we can solve it faster?

OpenStudy (skullpatrol):

I'm looking for common factors... Perhaps the prime factorization of 2008, 2009, 2010, 2011, 2012, and 2013 will show some pattern.

OpenStudy (kevin):

I got it. Tq very much!

OpenStudy (skullpatrol):

Please share the answer...

OpenStudy (skullpatrol):

all I could find was it's approximately equal to 40/9

OpenStudy (princeharryyy):

Was it really division? @Kevin

OpenStudy (princeharryyy):

I was on mobile and I thought it was a series of something ------- in between . LOL

OpenStudy (princeharryyy):

easiest way>> use logarithms to solve.

OpenStudy (kevin):

Sorry, I write the question wrong. The true question like this : 3^2008 (10^2013 + 5^2012 x 2^2011) ------------------------------------ 5^2012 (6^2010 + 3^2009 x 2^2008) So the answer would be like this : 3^2008 ((2 x 5)^2013 + 5^2012 x 2^2011) ---------------------------------------- = 5^2012 (2 x 3)^2010 + 3^2009 x 2^2008) 3^2008 x 2^2013 x 5^2013 + 3^2008 x 5^2012 x 2^2011 ------------------------------------------------------ = 5^2012 x 2^2010 x 3^2010 + 5^2012 x 3^2009 x 2^2008 3^2008 x 2^2008 x 5^2012 (2^5 x 5 + 2^3) ----------------------------------------- = 3^2008 x 2^2008 x 5^2012 (3^2 x 2^2 + 3) Until this, we know what is the next step :D Pls correct my answer if you find something wrong.

OpenStudy (kevin):

Yes, it's division @princeharryyy It's algebra contains division in it

OpenStudy (princeharryyy):

You are doing it okay. Cancel out the common terms.

OpenStudy (princeharryyy):

(2^5 x 5 + 2^3) -------------- = (32*5 +8)/(36+3) = 168/39 (3^2 x 2^2 + 3)

OpenStudy (princeharryyy):

168/39 = 56/13

OpenStudy (kevin):

yeah, that's my final answer :D Sorry If I write the question wrong and not clearly

OpenStudy (kevin):

Tq for your help

OpenStudy (princeharryyy):

It's okay. All I thought was that there were many terms. And the only possible question would have been to find the remainder like your earlier problem. For these kind of division problems you always take out the common terms. and if you can't find anything common you can find individual values using logarithms and try to find the answer.

OpenStudy (skullpatrol):

$$\large\dfrac{3^{2008} (10^{2013} + 5^{2012} 2^{2011})}{5^{2012} (6^{2010} + 3^{2009} 2^{2008})}$$ $$\large=\dfrac{3^{2008} ((2\cdot5)^{2013} + 5^{2012} 2^{2011})}{5^{2012} ((2\cdot3)^{2010} + 3^{2009} 2^{2008})}$$ $$\large=\dfrac{3^{2008} (2^{2013}\cdot5^{2013} + 5^{2012} 2^{2011})}{5^{2012} (2^{2010}\cdot3^{2010} + 3^{2009} 2^{2008})}$$ $$\large=\dfrac{3^{2008} 2^{2011} (5^{2013} 2^2 + 5^{2012} )}{5^{2012} 2^{2008} (2^{2} \cdot3^{2010} + 3^{2009} )}$$ $$\large=\dfrac{3^{2008} 2^{2011} 5^{2012} (5 \cdot 2^2 +1 )}{5^{2012} 2^{2008} 3^{2009} (2^{2} \cdot3 + 1 )}$$ $$\large=\dfrac{3^{2008} 2^{2011} \cancel{5^{2012}} (5 \cdot 2^2 +1 )}{\cancel{5^{2012}} 2^{2008} 3^{2009} (2^{2} \cdot3 + 1 )}$$ $$\large=\dfrac{3^{2008} 2^{2011} (5 \cdot 2^2 +1 )}{ 2^{2008} 3^{2009} (2^{2} \cdot3 + 1 )}$$ $$\large=\dfrac{ 2^{2011} (5 \cdot 2^2 +1 )}{ 2^{2008} 3 ( 2^{2} \cdot3 + 1 )}$$ $$\large=\dfrac{ 2^{3} (5 \cdot 2^2 +1 )}{ 3 ( 2^{2} \cdot3 + 1 )}$$ $$\large=\dfrac{ 8 (5 \cdot 4 +1 )}{ 3 ( 4 \cdot3 + 1 )}$$ $$\large=\dfrac{ 8 (21 )}{ 3 ( 13 )}=56/13$$

OpenStudy (kevin):

Yeah, that's it It's the correct question...

OpenStudy (skullpatrol):

Good job :-)